Presented below is the comparative income statement of Edison Company for 2018 and 2017.

Edison Company
Income Statement
For the Year Ended November 30
2018 2017
Sales $188,611 $123,848
Cost of Goods Sold $110,424 $64,407
Gross Profit $78,187 $59,441
Operating Expenses
Administrative Expense $21,184 $10,508
Selling Expense $27,225 $22,419
Total Operating Expenses $48,409 $32,927
Other Income and Expenses
Interest Expense $1,251 $867
Income Before Income Tax $28,527 $25,647
Income Tax $8,558 $7,694
Net Income (Loss) $19,969 $17,953

a) Calculate the following ratios for both years.

Do not enter dollar signs or commas in the input boxes.
Round your answers to 2 decimal places.

2018 2017
Net Profit Margin Answer Answer
Time Interest Earned Answer Answer

b) In which year does the company have a better performance with respect to the net profit margin calculated in part a)?

Year:
Answer #1

a) Calculate the following ratios for both years.

Do not enter dollar signs or commas in the input boxes.
Round your answers to 2 decimal places.

2018 2017
Net Profit Margin 19969/188611 = 10.59% 17953/123848 = 14.50%
Time Interest Earned (19969+1251+8558)/1251 = 23.80 (17953+7694+867)/867 = 30.58

b) In which year does the company have a better performance with respect to the net profit margin calculated in part a)?

2017 Year